# Evaluating the Diversity and Quality of LLM Generated Content

### v1

@@ Abstract
Recent work suggests that preference-tuning techniques—including Reinforcement Learning from Human Preferences (RLHF) methods like PPO and GRPO, as well as alternatives like DPO—reduce diversity, creating a dilemma given that such models are widely deployed in applications requiring diverse outputs. To address this, we introduce a framework for measuring effective semantic diversity—diversity among outputs that meet quality thresholds—which better reflects the practical utility of large language models (LLMs). Using open-ended tasks that require no human intervention, we find counterintuitive results: although preference-tuned models—especially those trained via RL—exhibit reduced lexical and syntactic diversity, they produce greater effective semantic diversity than SFT or base models, not from increasing diversity among high-quality outputs, but from generating more high-quality outputs overall. We discover that preference tuning reduces syntactic diversity while preserving semantic diversity—revealing a distinction between diversity in form and diversity in content that traditional metrics often overlook. Our analysis further shows that smaller models are consistently more parameter-efficient at generating unique content within a fixed sampling budget, offering insights into the relationship between model scaling and diversity. These findings have important implications for applications that require diverse yet high-quality outputs, from creative assistance to synthetic data generation.
@@ 1 Introduction
As large language models (LLMs) increasingly serve as tools for ideation, synthetic data generation, and creative assistance, their ability to produce diverse yet high-quality outputs has become critically important. While recent advances have dramatically improved the performance of these models, relatively little attention has been paid to systematically measuring and optimizing for this dual objective of quality and diversity.
Consider a user asking an LLM to generate story premises or a researcher using an LLM to create synthetic training data. In these scenarios, the utility of the model depends not only on generating a coherent set of responses but on producing a set of outputs that cover a meaningful range of ideas or training examples. This stands in contrast to traditional LLM evaluation paradigms that optimize for a single correct answer.
The challenge, however, lies in defining what constitutes meaningful diversity. Diversity without quality can be trivial to achieve: random tokens are maximally diverse but utterly useless. What we seek instead is diversity among outputs that meet a threshold of quality or acceptability. We term this effective semantic diversity and argue that it represents a more accurate measure of an LLM’s practical utility in open-ended generation tasks.
This distinction is particularly relevant in the current landscape of post-training LLMs, where preference-tuned models, through methods like Direct Preference Optimization (DPO), Proximal Policy Optimization (PPO), and Group Relative Policy Optimization (GRPO), are becoming standard and often yield high quality LLMs. While these methods have been successful in aligning models with human preferences, their impact on output diversity has been disputed. The dominant research narrative remains that preference tuning harms diversity. Moreover, a great deal of other open questions remain, such as the impact of model size on diversity and the ability to generate unique content.
In this paper, we introduce a principled framework for measuring this notion of high-quality diversity that:
@@
  • Allows for meaningful comparison across model families and training techniques
Figure 1: Diversity and Quality metrics when modulating the Temperature parameter across different models. We use the CodeBertScore (Zhou et al., 2023) model for neural cosine diversity.
In Figure 1 we show our implementation of effective semantic diversity for three Llama-3.1 family models when modulating the temperature parameter. We find that on its own, a diversity metric like neural cosine diversity is highest for the base LLM despite an extremely low proportion of its generations being valid (quality). Our implementation of effective semantic diversity penalizes excessively-low and high temperature values as well as models that greatly struggle to generate coherent content.
Using our framework, our experiments and analysis reveal several counterintuitive findings. We find that preference-tuning reduces semantic diversity relative to supervised fine tuning (SFT): but only if we restrict ourselves to the subset of high-quality generations, which generally gets larger with preference-tuning. As a result, preference-tuned models produce greater effective diversity than their SFT and base counterparts, where the increase in the proportion of high-quality outputs more than compensates for the reduction of diversity within high-quality generations. Nevertheless, we also observe a pattern where preference tuning, and especially reinforcement learning (RL) algorithms, reduce syntactic and lexical structures while preserving semantic diversity, suggesting an important distinction between diversity of form and diversity of content that linguistics has long recognized. Furthermore, we also find a pattern where RL post-training is associated with significantly less lexical and syntactic diversity relative to DPO post-training. Interestingly, this pattern differs from the effect of increasing model size, where we find that increases in effective diversity are not associated with a reduction in lexical or syntactic diversity. Additionally, in an attempt to understand which models are most efficient per parameter in generating unique programs, for example in generating unique synthetic-data examples, we generally find that going smaller, to as small as 500 million parameters, is preferable.
These insights come at a critical moment as the world increasingly looks to LLMs to assist in open-ended tasks that span ideation, synthetic data creation, and creative assistance. Our insights deepen the understanding of how post-training strategies meaningfully affect diversity, and our framework can be used to evaluate how post-training strategies of the future meaningfully affect diversity.
@@ 2 Background and Related Work
LLM Alignment with Preference Tuning. As LMs (Bengio et al., 2000; Radford et al., 2019) have become more powerful, work has been done to improve their instruction following ability and to mitigate the likelihood of generating undersirable content. RLHF with PPO (Schulman et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022) has emerged as a highly effective technique in aligning LLMs with human preferences. Subsequently, numerous alternative methods to PPO have been proposed, such as DPO (Rafailov et al., 2024), Rejection Sampling (Touvron et al., 2023), and GRPO (Shao et al., 2024). In numerous works, it has been reported that over-optimizing for the reward model eventually leads to incoherent or undesirable outputs (Ziegler et al., 2019; Stiennon et al., 2020; Korbak et al., 2022). It is suggested that an optimization strategy which places all probability on the highest reward outcome is optimal for this loss function, and will inevitably lead to distribution collapse (Lanchantin et al., 2025). In preference-tuning LLMs, the KL-divergence regularization has been instrumental to mitigating this effect as it allows the preference-tuned model to keep some of the distributional properties of the base LLM (Korbak et al., 2022). Therefore in-theory, KL-regularized RL should preserve some attributes of diversity present in base models.
@@
Approaches to Measuring Diversity. Due to the cost of human evaluation, popular methods for automatically measuring diversity fall into lexical or neural approaches. In earlier days when LMs often could struggle to generate rich content, lexical metrics were often used to measure diversity. These generally involve calculating summary statistics using $n$-grams, such as Distinct-N (Li et al., 2016; Du & Black, 2019) and Self-BLEU (a modified BLEU metric) (Zhu et al., 2018) and are still popular today (Guo et al., 2024; Shaib et al., 2024). As early foundation models for language representation, such as BERT (Devlin et al., 2019) became popular and used to model sentence similarity (Zhang et al., 2019; Reimers & Gurevych, 2019) and code similarity (Feng et al., 2020; Zhou et al., 2023), they were eventually proposed to measure diversity in natural language (Lai et al., 2020; Tevet & Berant, 2021; Stasaski & Hearst, 2022), but have not been applied for programming language diversity.
@@
While lexical and neural diversity metrics have been used for evaluating if one LLM is more diverse than another  (Kirk et al., 2023; Guo et al., 2024), to our knowledge, no work has actually evaluated if neural diversity metrics reflect diversity of effective semantic content in LLM-generated outputs. When LMs may run the gamut from tiny models which may often produce incoherent content, to large and powerful models with highly sophisticated abilities, different safety attributes, and even variations in “style," it is unclear if existing diversity metrics are capable of robustly reflecting meaningful semantic content across such a wide range of distributions. Models used for evaluating diversity like Sentence-BERT (Reimers & Gurevych, 2019) are generally trained and evaluated on human text; however, for diversity they’re expected to generalize to the most diverse sets of LLM-generated text which may be off-distribution. The closest attempt to evaluate lexical and neural diversity metrics is found in Tevet & Berant (2021) which evaluates if neural diversity metrics can reflect the temperature parameter used for sampling as well as diversity in human written content: for the latter it was found that neural models underperform human judgment.
@@
Diversity of LLM Content. Due to the high cost of human evaluation, most insights into the nature of novelty and diversity of LLMs are built upon linguistic and neural measures of diversity. Zhang et al. (2020) introduce the notion of a quality/diversity tradeoff in natural language generation in evaluating sampling algorithms like nucleus sampling (Holtzman et al., 2019) and temperature sampling. McCoy et al. (2023) investigate whether smaller LMs exhibit linguistic novelty by evaluating combinations of $n$-grams not present in the training corpus. As expectations rose with the advent of more chat-based LLMs like ChatGPT (Ouyang et al., 2022; OpenAI et al., 2023), the ability of these models to serve as creative assistants and propose diverse responses has come into question. It has been noted that the first release of ChatGPT was incapable generating diverse jokes (Jentzsch & Kersting, 2023). In a more recent study, Kirk et al. (2023) fine-tuned 7B AlpacaFarm (Dubois et al., 2024) models with SFT and PPO for neural text summarization, and found PPO reduces the lexical and neural diversity of summarized content. Padmakumar & He (2023) show that human-written essays assisted by an RLHF-tuned model are less diverse than those assisted by a base model when assessed using neural diversity measures. Guo et al. (2024) benchmark the lexical, syntactic, and neural diversity of 7B parameter language models, but do not probe whether or not preference-tuned models are more / less diverse diverse than SFT or base model alternatives. A separate line of work has probed LLMs from the perspective of social cognition to understand if LLMs reflect the same opinions and conceptual associations as macro-level populations of humans (Santurkar et al., 2023; Murthy et al., 2024). While these works have found that LLMs may not reflect the diversity of entire populations of people, it is highly unclear how this would impact the nature of effectively generating diverse content especially when placed under competing constraints of quality and helpfulness.
@@ 3.1 Problem Formulation
Validity function. To determine diversity at the individual prompt level, we define a binary validity function $V:\mathcal{G}\to\{0,1\}$, where $\mathcal{G}$ is the space of generations:
$V(g)=\begin{cases}1&\text{if }g\text{ is valid (i.e., meets quality threshold), }\\ 0&\text{otherwise.}\end{cases}$
@@
where $\mathrm{Set}(\cdot)$ denotes the set of unique elements and ||$\cdot$|| ⋅ | denotes cardinality. This measures the proportion of generations that are both valid/high-quality and semantically unique within the group of generations.
However, we found that Equation 2 can be confounded by the number of samples under consideration if analyzing a variable-sized subset of valid-only programs; see Section A.2 for more detail. To address potential sample size effects, we adopt the pairwise diversity metric suggested in  Tevet & Berant (2021):
@@
where the semantic distance function $d_{\text{sem}}:\mathcal{G}\times\mathcal{G}\to\{0,1\}$ is symmetric, i.e., $d_{\text{sem}}(g_{i}^{j},g_{i}^{k})=d_{\text{sem}}(g_{i}^{k},g_{i}^{j})$, and defined as:
@@
This pairwise approach normalizes by the total number of possible pairs, making it robust to variations in the number of valid generations across different prompts.
3.2 Dataset, Validity, and Semantic Equivalence Checking
@@
Function Signature: Write a function f(inputs) that processes the list of tuples where each tuple contains four real numbers. {minted}[fontsize=,frame=single,autogobble,breaklines]python from typing import List, Tuple def f(inputs: List[Tuple[float, float, float, float]]): ”’ inputs: a list of tuples, where each tuple contains four real numbers ”’
@@
We chose to instantiate methods of validity and semantic equivalence checking using programs. Where possible, we found it preferable to avoid using LLMs to evaluate LLMs for diversity, especially if we seek to evaluate increasingly powerful LLMs with weaker LLMs. And furthermore, we do this because there has been no empirical work studying the robustness of this general approach for comparing LLM content. Research in programming language semantics has a rich history in the formalization of program correctness, termination, and semantic equivalence (Hoare, 1969; Plotkin, 1981; Floyd, 1993; Pierce, 2002). Using program execution on test cases, we can confirm two programs are semantically distinct if they compute different values over the same suite of input test cases [^1] N.B. Test cases may fail to test edge cases where two programs may differ. However, instead of over-reporting correctness, this phenomenon would only penalize diversity when two programs are equivalent on most test cases and hypothetically differ on some edge-case, which we find tolerable.. Additionally, we can impose constraints of validity or quality that are easy to check: such as the program always returning a value without syntax or other runtime errors.
@@
We constructed our dataset by adapting competitive programming-style problems into open-ended abstracted programming tasks. In Figure 2, we show an example problem description from our dataset. For each problem description in our dataset, we provide an “Input Description” in natural language specifying the input format, an “Example Input” demonstrating potential inputs the function would handle, and a “Function Signature” providing a concrete specification of the function name and typing hints for the inputs. We manually reviewed, and if necessary, edited all final descriptions, removing any potential references to the original programming task, standardizing the function name to $f(...)$, and using highly generic argument names. We used competitive programming problems from CodeNet (Puri et al., 2021) and accompanying test cases from AlphaCode (Li et al., 2022) as a starting point for our dataset. In total we have 108 unique programming tasks / prompts that were adapted from CodeNet into this open-ended format. Using each of these, we can sample an arbitrary number of generations to calculate $\text{Div}_{m}(\mathcal{P}_{i})$. Further details about the test set, its creation, and our execution environment are provided in Section A.3.

### v2

# arXiv:2504.12522v2[cs.CL] 26 Feb 2026

@@ Abstract
Recent work suggests that preference-tuning techniques—such as Reinforcement Learning from Human Preferences (RLHF) methods like PPO and GRPO, as well as alternatives like DPO—reduce diversity, creating a dilemma given that these models are widely deployed in applications requiring varied outputs. We argue that diversity without consideration of quality has limited practical value. To address this issue, we introduce a framework for measuring effective semantic diversity—diversity among outputs that meet quality thresholds—which better reflects the practical utility of large language models (LLMs). Using open-ended tasks that require no human intervention, we find counterintuitive results: when using diversity metrics that do not explicitly consider quality, preference-tuned models—particularly those trained via RL—often produce outputs with lower diversity; however, these same preference-tuned models generate greater effective semantic diversity than SFT or base models. Our analysis further shows another trend: while larger models may exhibit greater effective semantic diversity than smaller models, the smaller models are consistently more parameter-efficient at producing unique content within a fixed sampling budget. These findings have practical implications for applications that require diverse yet high-quality outputs, from creative assistance to synthetic data generation.
@@ 1 Introduction
As large language models (LLMs) increasingly serve as tools for ideation, synthetic data generation, and creative assistance, their ability to produce diverse yet high-quality outputs has become critically important. While recent advances have dramatically improved model performance, relatively little attention has been paid to systematically measuring and optimizing for the dual objective of quality and diversity.
Consider a user asking an LLM to generate story premises or a researcher using an LLM to create synthetic training data. In these scenarios, the utility of the model depends not only on producing a coherent set of responses but also on generating outputs that span a meaningful range of ideas or examples. This contrasts with traditional LLM evaluation paradigms, which often optimize for a single correct answer.
The challenge lies in defining what constitutes meaningful diversity. Diversity without quality is trivial to achieve—random tokens are maximally diverse but utterly useless. What is needed instead is diversity among outputs that meet a threshold of quality or acceptability. We refer to this as effective semantic diversity and argue that it is a more accurate measure of an LLM’s practical utility in open-ended generation tasks.
This distinction is particularly relevant in today’s landscape of post-training LLMs, where preference-tuned models—trained via methods such as Direct Preference Optimization (DPO), Proximal Policy Optimization (PPO), and Group Relative Policy Optimization (GRPO)—are increasingly the standard. These methods have been highly successful at aligning models with human preferences, yet their impact on output diversity remains disputed. The dominant research narrative is that preference tuning harms diversity, and many open questions persist, including the effect of model size and the ability to generate unique content.
We introduce a principled framework for measuring high-quality diversity that:
@@
  • Enables meaningful comparison across model families and training techniques
Figure 1: Diversity and Quality metrics on our open-ended programming dataset when modulating the Temperature parameter across different models. We use the CodeBertScore (Zhou et al., 2023) model for neural cosine diversity.
In Figure˜1 we illustrate our implementation of effective semantic diversity for three Llama-3.1 family models under varying temperature settings. Using neural cosine diversity alone, the base LLM appears most diverse, despite producing an extremely low proportion of valid (high-quality) generations. Our effective semantic diversity metric penalizes both excessively low and excessively high temperatures, as well as models that struggle to generate coherent content.
With this framework, our experiments reveal insights into preference-tuning and the diversity-quality trade-off. Using diversity metrics that do not explicitly consider quality, preference tuning may indeed reduce diversity relative to supervised fine-tuning (SFT). However, in both code generation and creative writing, we find that preference-tuned models exhibit greater effective semantic diversity. A deeper analysis in code generation shows that when restricting to high-quality generations only, preference-tuned models produce less semantically diverse content within that subset. Nonetheless, preference tuning increases the proportion of high-quality outputs to such an extent that this gain outweighs the reduction in diversity per high-quality sample.
In code generation, preference tuning—especially reinforcement learning (RL)—is associated with reduced lexical and syntactic diversity but no loss in semantic diversity. For open-ended creative writing, preference tuning is linked to greater diversity in lexical patterns. Finally, when evaluating parameter efficiency for generating unique programs within a fixed sampling budget—for example, when creating unique synthetic data—we find that smaller models, down to around 500 million parameters, are often the most efficient choice.
@@ 2 Background and Related Work
LLM Alignment with Preference Tuning. As LMs (Bengio et al., 2000; Radford et al., 2019) have become more powerful, work has been done to improve their instruction-following ability and to mitigate the likelihood of generating undesirable content. RLHF with PPO (Schulman et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022) has emerged as a highly effective technique for aligning LLMs with human preferences. Subsequently, numerous alternative methods to PPO have been proposed, such as DPO (Rafailov et al., 2024), Rejection Sampling (Touvron et al., 2023), and GRPO (Shao et al., 2024). In numerous works, it has been reported that over-optimizing for the reward model eventually leads to incoherent or undesirable outputs (Ziegler et al., 2019; Stiennon et al., 2020; Korbak et al., 2022). It is suggested that an optimization strategy that places all probability on the highest-reward outcome is optimal for this loss function but will inevitably lead to distribution collapse (Lanchantin et al., 2025). In preference-tuning LLMs, KL-divergence regularization has been instrumental in mitigating this effect, as it allows the preference-tuned model to retain some of the distributional properties of the base LLM (Korbak et al., 2022). Therefore, in theory, KL-regularized RL should preserve some attributes of diversity present in base models.
@@
Approaches to Measuring Diversity. Due to the high cost of human evaluation, popular methods for automatically measuring diversity typically fall into either lexical or neural approaches. In the earlier days, when LMs often struggled to generate rich content, lexical metrics were commonly used to assess diversity. These metrics generally involve calculating summary statistics over $n$-grams, such as Distinct-N (Li et al., 2016; Du & Black, 2019) and Self-BLEU (a modified BLEU metric) (Zhu et al., 2018), and remain widely used today (Guo et al., 2024; Shaib et al., 2024). As early foundation models for language representation, such as BERT (Devlin et al., 2019), gained popularity and were adopted for modeling sentence similarity (Zhang et al., 2019; Reimers & Gurevych, 2019) and code similarity (Feng et al., 2020; Zhou et al., 2023), they were eventually proposed for measuring diversity in natural language (Lai et al., 2020; Tevet & Berant, 2021; Stasaski & Hearst, 2022). However, they have not yet been applied to measuring diversity in programming languages.
@@
While lexical and neural diversity metrics have been used to evaluate whether one LLM is more diverse than another (Kirk et al., 2023; Guo et al., 2024), to our knowledge no prior work has assessed whether neural diversity metrics truly capture the diversity of effective semantic content in LLM-generated outputs. Given that LMs range from small models that often produce incoherent content to large, powerful models with sophisticated capabilities, varying safety attributes, and distinctive “styles," it remains unclear whether existing diversity metrics can robustly reflect meaningful semantic content across such a wide range of distributions. Models used for evaluating diversity, such as Sentence-BERT (Reimers & Gurevych, 2019), are typically trained and evaluated on human-authored text; for diversity assessment, however, they are expected to generalize to highly diverse, potentially off-distribution sets of LLM-generated text. The closest prior attempt to evaluate lexical and neural diversity metrics is Tevet & Berant (2021), which examined whether neural diversity metrics can capture variation induced by the sampling temperature parameter and diversity in human-written content; in the latter case, neural models were found to underperform human judgment.
@@
Diversity of LLM Content. Given the expense of human evaluation, most insights into the novelty and diversity of LLM outputs are based on linguistic and neural measures of diversity. Zhang et al. (2020) introduce the notion of a quality-diversity trade-off in natural language generation when evaluating sampling algorithms such as nucleus sampling (Holtzman et al., 2019) and temperature sampling. McCoy et al. (2023) investigate whether smaller LMs exhibit linguistic novelty by evaluating combinations of $n$-grams absent from the training corpus. With the advent of more chat-based LLMs like ChatGPT (Ouyang et al., 2022; OpenAI et al., 2023), expectations for these models to serve as creative assistants and produce diverse responses have risen, but their capabilities have been questioned. For example, the first release of ChatGPT was reported to be incapable of generating diverse jokes (Jentzsch & Kersting, 2023). In a more recent study, Kirk et al. (2023) fine-tune 7B AlpacaFarm (Dubois et al., 2024) models with SFT and PPO for neural text summarization and find that PPO reduces both lexical and neural diversity in the summaries. Padmakumar & He (2023) show that human-written essays assisted by an RLHF-tuned model are less diverse than those assisted by a base model when assessed using neural diversity measures. Guo et al. (2024) benchmark the lexical, syntactic, and neural diversity of 7B parameter language models but do not probe whether preference-tuned models are more or less diverse than SFT or base alternatives. A separate line of work has examined LLMs from a social cognition perspective, investigating whether LLMs reflect the opinions and conceptual associations of macro-level human populations (Santurkar et al., 2023; Murthy et al., 2024). Recent work has also explored the role of diversity metrics in evaluating and improving synthetic data quality for LLMs: Yu et al. (2023) investigate LLMs as attributed training data generators; Divekar & Durrett (2024) propose SynthesizRR, a retrieval-augmented framework for generating synthetic datasets with greater lexical and semantic diversity; Chen et al. (2024) study the downstream impact of synthetic data diversity on training LLMs; and Miranda et al. (2025) introduce a quantitative metric, the diversity coefficient, to measure variability in natural language datasets. While these works find that LLMs may not reflect the diversity of entire human populations, it remains unclear how this impacts their ability to generate effectively diverse content—particularly when faced with competing constraints of quality and helpfulness.
@@ 3.1 Problem Formulation
Validity function. To determine diversity at the individual prompt level, we define a validity function $V:\mathcal{G}\to[0,1]$, where $\mathcal{G}$ is the space of generations.
@@
where $\mathrm{Set}(\cdot)$ denotes the set of unique elements and $|\cdot|$ denotes cardinality. This measures the proportion of generations that are both valid/high-quality and semantically unique within the group of generations.
However, we found that Equation˜2 can be confounded by the number of samples under consideration if analyzing a variable-sized subset of valid-only programs; see Section˜A.2 for more detail. To address these concerns, we adopt the pairwise diversity metric suggested in  Tevet & Berant (2021):
@@
where the effective semantic distance function $d_{\text{sem}}:\mathcal{G}\times\mathcal{G}\to\{0,1\}$ is defined as:
@@
This pairwise approach normalizes by the total number of possible pairs, making it robust to variations in the number of valid generations across different prompts. However, in domains such as natural language where strict notions of semantics do not apply, we define a soft approach to measuring $d_{\text{sem}}$ as:
$d_{\text{sem}}(g_{i}^{j},g_{i}^{k})=V(g_{i}^{j})\times V(g_{i}^{k})\times(1-Sim(g_{i}^{j},g_{i}^{k}))$ (5)
Here $Sim:\mathcal{G}\times\mathcal{G}\to[0,1]$ is a similarity function that does not explicitly consider quality, for example, assigned by a human or AI judge.
3.2 Dataset, Validity, and Semantic Equivalence Checking for Programs
@@
.................................................................................................................................... Function Signature: Write a function f(inputs) that processes the list of tuples where each tuple contains four real numbers. from typing import List, Tuple def f(inputs: List[Tuple[float, float, float, float]]): ”’ inputs: a list of tuples, where each tuple contains four real numbers ”’
@@
We chose to instantiate methods of validity and semantic equivalence checking using both programs and natural language. We found it advantageous to use programs to avoid the scenario of using LLMs to evaluate LLMs for diversity, especially if we seek to evaluate increasingly powerful LLMs with weaker LLMs. Furthermore, there has been relatively little empirical work that studies the robustness of using LLMs to judge the diversity of LLM-generated content. Lastly, our implementation of effective semantic diversity for programs avoided the costs of utilizing commercial LLMs as a judge. Research in programming language semantics has a rich history in the formalization of program correctness, termination, and semantic equivalence (Hoare, 1969; Plotkin, 1981; Floyd, 1993; Pierce, 2002). Using program execution on test cases, we can confirm two programs are semantically distinct if they compute different values over the same suite of input test cases [^1] N.B. Test cases may fail to test edge cases where two programs may differ. However, instead of over-reporting correctness, this phenomenon would only penalize diversity when two programs are equivalent on most test cases and hypothetically differ on some edge-case, which we find tolerable.. Additionally, we can impose constraints of validity or quality that are easy to check: such as the program always returning a value without syntax or other runtime errors.
@@
We constructed our dataset by adapting competitive programming-style problems into open-ended abstracted programming tasks. In Figure˜2, we show an example problem description from our dataset. For each problem description in our dataset, we provide an “Input Description” in natural language specifying the input format, an “Example Input” demonstrating potential inputs the function would handle, and a “Function Signature” providing a concrete specification of the function name and typing hints for the inputs. We manually reviewed, and if necessary, edited all final descriptions, removing any potential references to the original programming task, standardizing the function name to f()f(...), and using highly generic argument names. We used competitive programming problems from CodeNet (Puri et al., 2021) and accompanying test cases from AlphaCode (Li et al., 2022) as a starting point for our dataset. In total we have 108 unique programming tasks/prompts that were adapted from CodeNet into this open-ended format. Using each of these, we can sample an arbitrary number of generations to calculate $\text{Div}_{m}(\mathcal{P}_{i})$. Further details about the test set, its creation, and our execution environment are provided in Section˜A.3.
3.3 Dataset for Natural Language
For natural language, we created our dataset by taking a subsample of 100 creative writing prompts from the WritingPrompts dataset (Fan et al., 2018). We took a random subset of the test set, prioritized [WP] tagged prompts, and avoided prompts that could generate sensitive or inappropriate stories.